A helicopter flew 180 miles 1.5 hours with a tail wind. The return trip took 2 hours with a head wind. What was the speed of the helicopter in still air? What was the speed of the wind?
Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

180/1.5 = 120 mph
180/2 = 90 mph
The airspeed is the average (120+90)/2 = 105 mph
---------
The windspeed is the difference = 15 mph

Answer by ewatrrr(24785)


Hi,
*Note: distance= rate * time Or rate = distance/time
rate with the wind = 180/1.5 = 120mph
rate against the wind = 180/2 = 90mph
s + w = 120mph
s - w = 90mph
2s = 210mph
s = 105mph, the speed of the helicopter
wind = 15mph